solve for y:

-2y^2 + 7y = -15

Answer by checkley77(12844)

-2y^2+7y=-15
-2y^2+7y+15=0
(-2y-3)(y-5)=0
-2y-3=0
-2y=3
y=-3/2 answer.
y-5=0
y=5 answer.